OpenStudy (anonymous):

Jackson has a previous balance of $767 on a credit card with a 16.5% APR compounded monthly. If he made a payment of $49 this month, what is the new balance on his credit card?

OpenStudy (anonymous):

okay so i need to add all the interest rates and charges? then subtract them form 767?

OpenStudy (anonymous):

but that about the 49?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

why do i need to know the apy i thought i needed to find the balance

OpenStudy (anonymous):

so 85.56?

OpenStudy (anonymous):

Aiko, your responses seem to be using different methods! After the first reply you seemed to go on to explain a different method. I believe \[(1+\frac{apr}{n})^n = apy\] is the correct equation if the effective APR is given.

OpenStudy (anonymous):

(where n is the number of times interest is added per year, apr is the nominal APR, and apy is the effective APR)

OpenStudy (anonymous):

okay so is 85.56 the correct APY? and if so what do i do with that to get the balance?

OpenStudy (anonymous):

add and subtract what? that makes no sense to me, i dont know what any of those have to do with APY. or what to do with the previous balance and the payment of 49

OpenStudy (anonymous):

The question's wording is strange. My assumption is that the money is paid out and then interest is calculated? This would make the new balance \[(767-49)(1+\frac{apr}{n})\] If this is indeed what the question asks, the answer can be calculated using the equation in my previous post, remembering that n=12 since the balance is done monthly.

OpenStudy (anonymous):

ok i think its better if @Indivicivet explains and i stay out becuase i am producing to much confusion for u @OtakuPrincess

OpenStudy (anonymous):

okay i got 1705.25 but thats not one of my choices :/

OpenStudy (anonymous):

oh wait i think i forgot to move the decimal! 727.87?

OpenStudy (anonymous):

thats a lot closer to one of my choices

OpenStudy (anonymous):

My mistake, in the first equation it should be 1+apy instead of apy. Rearranging the first equation I posted gives: \[1+\frac{apr}{n} = \sqrt[n]{1+apy}\] n=12 and apy = 0.165, therefore: \[1+\frac{apr}{n}\approx1.013\] Inserting into the second equation gives \[Balance\approx(767-49)(1.013) = 718*1.013\approx727.196\] I used the exact figures throughout my calculation. What are your choices? I might've made a mistake but otherwise, if they don't match up with this, I'd expect the question may be asking something different. As I said before, it was badly worded.

OpenStudy (anonymous):

$718.00 $728.55 $844.56 $893.56 OOHH okay thanks i get that and its super close to 728.55

OpenStudy (anonymous):

Wait, $718 is an answer? I feel like this could be a trick question based on awful wording.